OpenStudy (anonymous):

how can i derivate the following derivative? [ (√25 + x^2)/2] + [(6-x)/4]

OpenStudy (saifoo.khan):

\[\frac{\sqrt{25+x^2}}{2}+ \frac{6-x}{4}\]Like this?

OpenStudy (anonymous):

\[\frac{ \sqrt{25 + x ^{2}} }{ 2 } + \frac{ 6-x }{ 4 }\]

OpenStudy (anonymous):

yeah!

OpenStudy (saifoo.khan):

I'll do it this way: \[\Large \frac{2\sqrt{25+x^2}+6-x}{4} \]

OpenStudy (saifoo.khan):

Now: \[\Large \frac{1}{4} \times [\frac{d}{dx} 2(25+x^2)^{0.5}+6-x]\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

but how did u get the \[\frac{ 1 }{ 4 }\] ??

OpenStudy (saifoo.khan):

Did you understand the first step? I took LCM in it. In next step, since constants stay the same, i took it out.

OpenStudy (saifoo.khan):

You can write my second step as: \[\Large [2\sqrt{25+x^2}+6-x] \times [\frac{1}{4}]\]

OpenStudy (anonymous):

ooh, i understood, ill try to do the following,,, :) thank you very much

OpenStudy (saifoo.khan):

You're welcome. :)

OpenStudy (saifoo.khan):

Sure. Where are you stuck?

OpenStudy (anonymous):

uhmm once im trying to solve the derivative i get stuck!, i get something like this\[\frac{ 1 }{ 4\sqrt{25+x ^{2}} }-\frac{ 1 }{ 4 }\] is that right?

OpenStudy (saifoo.khan):

You should get: \[\frac{x}{2\sqrt{x^2+25}}- \frac14\]

OpenStudy (anonymous):

uhm how did u get that? :$

OpenStudy (saifoo.khan):

Let me show it to you.

OpenStudy (saifoo.khan):

I'll attach a photo.

OpenStudy (anonymous):

yepp, thank you so!

OpenStudy (anonymous):

thank you so much!! uhmm but i got a question, why you get 0.5 in the denominator?

OpenStudy (saifoo.khan):

0.5 = 1/2 same thing.

OpenStudy (saifoo.khan):

Like if i multiply "a" with 0.5, i will get 0.5a or 1/2 a